an airplane A leaves los angeles for denver at a speed of 450 mph. thirty minutes later a plane B going from denver to los angeles leaves denver which is 851 miles from los angeles at a speed of 489 mph. when they meet how far are they apart from denver?
Answer by stanbon(75887)

an airplane A leaves los angeles for denver at a speed of 450 mph. thirty minutes later a plane B going from denver to los angeles leaves denver which is 851 miles from los angeles at a speed of 489 mph. when they meet how far are they apart from denver?
-----
LA to Denver DATA:
rate = 450 mph ; time = x hrs ; distance = r*t = 450x miles
-----
Denver to LA DATA:
rate = 489 mph ; time = x-0.5 hrs ; distance = r*t = 489(x-0.5) miles
--------
Equation:
dist + dist = 851 miles
450x + 489x - 489(0.5) = 851
939 x = 851 + 244.5
-----
939x = 1095.5
x = 7/6 hr.
---
Distance from Denver: (7/6)489 = 570.5 miles
==================================================
Cheers,
Stan H.
====================